What is the change in internal energy (in J) of a system that absorbs 0.297 kJ of heat from its surroundings and has 0.750 kcal of work done on it? Give your answer in scientific notation. × 10,
Answer #1

According to the first law of thermodynamics, the change in internal energy (\DeltaE) is the sum of heat (q) and work (w):

\DeltaE = q + w

Since heat is absorbed by the system from the surroundings, heat (q) is positive; also, work (w) is positive since work is done on the system.

Converting the values of heat to work to Joules, we get:

q = 0.297 kJ \times 103 J \div 1 kJ [1 kJ = 103 J]

= 297 J

w = 0.750 kcal \times 103cal \div 1 kcal \times 4.184J \div 1 cal [1 kcal = 103cal ; 1 cal = 4.184 J]

= 3138 J

Therefore, \DeltaEint = q + w = +297 J - [-3138 J] = 3435 J

There is an increase in internal energy.

3435 J can be written as :

Change in internal energy = 3.435 \times 103 J